find two numbers such that the larger is 1 less than 3 times the smaller. three times the larger is 5 more than 8 times the smaller.
Answer by elima(1433)

Let x = the smaller number
3x-1=the larger number
three times the larger is 5 more than 8 times the smaller.
3(3x-1)=5+8x
9x-3=5+8x
9x-8x=5+3
x=8
The smaller number = 8
the larger number= 3(8)-1=24-1=23
larger number = 23
=)